# General relativity and the double-zero eigenvalue

### v1

@@ Abstract
For a given nonlinear system of dynamical equations, the set of all stable perturbations of the system, the ‘versal unfolding’, is fully determined by bifurcation theory. This leads to a complete depiction of all possible qualitatively different solutions and their metamorphoses into new topological forms, the versal families being dependent on a suitable number - the codimension - of independent parameters. In this paper, we introduce and elaborate on the idea that gravitational phenomena are bifurcation problems of finite codimension. We deploy gravitational bifurcations for five standard problems in general relativity, namely, the Friedmann models, the Oppenheimer-Snyder black hole, the issue of a focussing state associated with the evolution of congruences of causal geodesics in the context of cosmology and black holes, the crease flow on black hole event horizons, and the Friedmann-Lemaître equations. We also comment on more general emerging aspects of the application of bifurcation theory to problems involving the gravitational interaction, and the possible relations between gravitational versal unfoldings and those of the Maxwell, Dirac, and Schrödinger equations.
1 Introduction
There is a somewhat concealed feature of gravitational field equations and their possible reductions in specialized situations describing phenomena of interest in cosmology, black holes, and other areas in gravity research. It is shared for example by the Einstein equations, or by analogous ones in modified gravity or string theories, and also by ‘optical’ type systems describing the evolution of congruences of causal curves or higher-dimensional gravitating objects in spacetime. This feature can be well described both physically and mathematically by the statement that:
@@ 1 Introduction
The main message of this paper is that gravitation is a structurally unstable phenomenon and must be treated using an approach where this feature does not disappear altogether in the process but persists. In particular, the observed cosmic structures could have never been formed and persisted had gravity been structurally stable. How then does gravitational instability in its various forms manifest itself as structural instability of the governing gravitational equations? How can one describe the latter in a mathematically precise yet physically consistent and plausible way? What are the main physical implications predicted by gravitational structural instability alone that could never have arisen had we treated the equations of a particular gravitational setup as being structurally stable?
Physically speaking, bifurcation theory may be defined as the study of structurally unstable (also called ‘dispersive’) phenomena, and while its precise nature as a mathematical theory is probably unfamiliar to, or perhaps even might be considered as peripheral at best by, experts in the field of gravitation, the purpose of this paper is rather to suggest that there is a fundamental link between gravity and bifurcation theory and explore this relation using methods of the singularity theory of functions [^1] For further discussion and proofs of the general bifurcation theory statements used below the reader is invited to consult Refs. [1]-[18], while for all statements related to gravitational bifurcations, we refer to [19]-[22]..
Structural stability [^2] We shall use the words structural stability, hyperbolicity, and non-degeneracy as synonymous, and the word bifurcation as synonymous to structural instability, dispersive behaviour, non-hyperbolicity, and degeneracy. We shall also generally use ‘bifurcation theory’ as a generic phrase describing all three mathematical theories of ‘bifurcations’, ‘singularities’ (of functions), and ‘catastrophes’. is a mathematically well-known concept of stability that refers to the stability of the laws themselves, that is to the dynamical equations governing the phenomena, as opposed to the stability of their specific solutions. However, it imposes very strong restrictions to the solutions describing different manifestations of the physical phenomena associated with these laws. For a system described by such equations, the spectrum of possible behaviours is restricted by the condition that any nearby system, i.e., any small perturbation of the system, behaves exactly as the original one [23]-[25].
This demand of structural stability on a given system leads to only three possible behaviours for any solution: stable, unstable, or saddle-like. That is, either there is nothing new, i.e., a solution rapidly dies out and the system returns to its unperturbed state, or diverges, ripping the system apart, or - as perhaps the worst alternative - a mixture of the two that results in an unstable saddle-like state. Consequently, such a system is in some sense always ‘trapped in itself’ being unable to pass to another, qualitatively different, state like for example one of those we observe around us (this will be made precise below).
There is a great difference between structurally stable and dispersive systems of dynamical equations [26], [27]. For our present purposes, a very characteristic manifestation of that difference is provided by the phenomenon of gravitational instability, as in the problem of cosmological structure formation where, albeit usually treated in the linear approximation [28]-[38], the universe was very smooth but has become very inhomogeneous on small scales as cosmological observations clearly suggest. The universe ‘settles down’ to a novel state which albeit a changing one does not suddenly disappear but ‘persists’ for a time comparable to the observed lifetime of the system.
In addition to the above, gravitational systems share two further properties which are in fact very akin to those of typical structurally unstable systems:
@@
    They possess parameters whose values are not really known exactly.
@@
    They are described by differential equations which are very sensitive to even very small changes or modifications of their defining terms.
@@
The fluid parameter $\gamma$, or a cosmological constant are examples of the first property, while any randomly picked system of equations describing a gravity problem may be indicative of the second. Although a demand of structural stability is understandable as it implies that a property possessed by such a system is then physically very plausible, such a restriction imposed on an unstable system satisfying the two above conditions may lead to non-characteristic results attributed to the original equations and leave out other behaviours which might lie closer to the nature of the system’s instabilities.
@@
Another related issue is what one really means by ‘stability’ in the unstable context of a structurally unstable system? How may one classify the various unstable modes available in a given gravitational problem described and governed by a structurally unstable set of dynamical equations? Such issues point to a further consideration of related questions. For example, how can one perturb a structurally unstable system? What is the possible range of such perturbations, and how are the solutions of the original system are being affected by the latter? What is the meaning of the phrase ‘small perturbation’ for a structurally unstable situation where there may be almost ignorable perturbations that have the ability to completely corrupt all orbits?
@@
We shall return to these questions in the coming sections of this paper, where we shall introduce the approach of bifurcation theory to this problem. Basically, very roughly speaking, this approach to the problem of general gravitational instabilities dictates that: one must first recognise the kinds of degeneracy of the original system by bringing it in a ‘normal form’; secondly, embed the system of dynamical equations in a suitable larger parametric family customized to the given problem; and thirdly, prove that this family contains generically all possible stable perturbations of the original system. This is a long path but eventually becomes very rewarding. One acquires control over all degeneracies of the original system, now in a structurally stable framework where each unstable aspect persists, and one is in this way led to a spectrum of all possible behaviours associated with the original system.
@@
In Section 2, we discuss the main differences between hyperbolic and nonhyperbolic systems, and in Section 3 we formulate the main problem addressed in this work and write down the central objectives in the search of gravitational bifurcations. Section 4 gives a brief exposition of the main methods of bifurcation theory. In Section 5, we study the picture which emerges from an application of these methods to some of the most basic and well-studied gravity problems in general relativity, cosmology, and black holes. In the last Section, we discuss more general issues which emerge from gravitational bifurcations.
@@
This work is based on the results of Refs. [19]-[22], however, we skip all details of the long arguments and calculations needed to prove many of the statements made in those references and referred to here. After reading the present paper, we hope that the reader will be sufficiently motivated to at least look at some of the proofs in those references.
2 Hyperbolic vs. dispersive systems
2.1 General comments
@@
Besides in the theory of small fluctuations mentioned above, an essentially ‘hyperbolic’ approach based on the concept of structural stability is as a rule the one used in other dynamical treatments of gravity problems, cf. e.g., [39]-[41] and refs. therein, and has its roots in the standard theory of (hyperbolic) dynamical systems [23]-[25]. As is well-known this practice leads to systems with (usually complicated) hyperbolic equilibria, and so imposes two very important conditions on the gravity systems studied in this way: firstly, near such equilibria they become topologically equivalent to their linear parts, and secondly, ‘nearby’ nonlinear systems behave like their unperturbed counterparts.
We shall consider structurally unstable systems only and not follow the more or less common approach of restricting the parameters present in a gravity problem to suitable intervals, and then treat a possibly dispersive system as being structurally stable. Our approach has a very high applicability not only in cosmology but in other areas as well (for a description of possible cosmological areas of application of these ideas, cf. e.g., [42]).
@@
To motivate our approach to gravitational bifurcation theory employed in following Sections, let us recall the difference between a hyperbolic and a non-hyperbolic system (a hyperbolic system is one whose all equilibria are hyperbolic). This difference arises in two ways: while equilibria in a hyperbolic system are always isolated and the behaviour of solutions near them is qualitatively the same as that of their linear parts, dispersive systems exhibit bifurcating behaviour and as a rule comprise multiple solutions intertwined with dramatic changes in the behaviour, while their global dynamics is restricted to only those dimensions which correspond to their non-hyperbolic eigenvalues.
2.2 Hyperbolic dynamics

### v2

# Structural stability and general relativity

@@ Abstract
We review recent developments in structural stability as applied to key topics in general relativity. For a nonlinear dynamical system arising from the Einstein equations by a symmetry reduction, bifurcation theory fully characterizes the set of all stable perturbations of the system, known as the ‘versal unfolding’. This construction yields a comprehensive classification of qualitatively distinct solutions and their metamorphoses into new topological forms, parametrized by the codimension of the bifurcation in each case. We illustrate these ideas through bifurcations in the simplest Friedmann models, the Oppenheimer-Snyder black hole, the evolution of causal geodesic congruences in cosmology and black-hole spacetimes, crease flow on event horizons, and the Friedmann-Lemaître equations. Finally, we list open problems and briefly discuss emerging aspects such as partial differential equation stability of versal families, the general relativity landscape, and potential connections between gravitational versal unfoldings and those of the Maxwell, Dirac, and Schrödinger equations.
“… versality implies the structural stability of the deformation.”
— V. I . Arnold (1994)
“In order that stationary processes may exist in a real system for a long time, they must be stable not only with respect to the coordinates and the velocities, but also with respect to small variations of the form itself of the differential equations describing the system … these [small variations of the differential equations] shall be assumed at first to be such as not to vary the order of the initial differential equation … such as not to force us to reject the idealization connected with the number of degrees of freedom. … Systems which are such as not to vary in their essential features for a small variation of the form of the differential equations we shall call “coarse” systems… .”
— A. A. Andronov, A.A. Vitt, and S. E. Khaikin (1966)
“There seems to be a time scale in all natural processes beyond which structural stability and calculability become incompatible …, and today the physicist sacrifices structural stability for computability.”
— R. Thom (1972)
“When does the phase portrait itself persist under perturbations of the vector field? This is the problem of structural stability.”
— M. W. Hirsch and S. Smale (1974)
“The word bifurcation … is used in a wide sense to indicate every qualitative topological metamorphosis of a picture under the variation of parameters on which the object being studied depends.”
— V. I . Arnold (1988)
1 Introduction and overview
1.1 General relativity and symmetry reduction
General relativity occupies a central position in modern physical theories. It naturally interfaces with other areas of theoretical physics, partially ‘unifying’ them via the Einstein equations: the stress tensor on the right-hand-side couples directly to the geometric (gravitational) sector. This coupling yields a rich array of physical predictions and profound effects many of which have been confirmed observationally. Whenever an exact solution of the Einstein equations is available [1], one performs a symmetry reduction that converts the Einstein partial differential equation (PDE) into a finite-dimensional system of ordinary differential equations (ODEs). Such reductions — common in homogeneous cosmology, black-hole problems, relativistic astrophysics, and gravitational-wave studies — ‘collapse’ the PDE to a lower–dimensional dynamical system (cf. e.g., [2]-[7]). Although some information is lost, the resulting ODE systems remain of significant mathematical and physical interest.
1.2 Analogs in reaction-diffusion and travelling waves
This symmetry-reduction-related situation is not unique to general relativity, but appears in many important — and unrelated both physically and mathematically to gravity — PDEs, notably reaction-diffusion equations, where one derives a ‘travelling-wave’ equation near the wave front. Such travelling wave equations are the exact analogues of the symmetry-reduced equations of general relativity, and while the properties of the reduced PDE cannot capture the essence of the full ‘unreduced’ PDE, most of the reduced equations obtained this way have a common property: they are structurally unstable. This creates an equally amazing variety of phenomena at the reduced-ODE level, and frame the fundamental question of the ODE-PDE correspondence (see [8]-[14] for reaction-diffusion and other types of equations). A main reason for the structural instability of travelling-wave-like equations is their degenerate nature and, as we shall demonstrate below, the same holds in general relativity. In particular, while the techniques to obtain stability at the PDE level for reaction-diffusion and the Einstein equations are very different, the symmetry-reduced dynamical systems typically obtained by such a process in the case of general relativity are typically structurally unstable as well. This instability has important implications for gravitational models, yet remains underexplored in the literature.
1.3 Structural stability vs. instability
Structural stability is a well-established mathematical concept concerning the robustness of dynamical laws under small perturbations of their defining equations, rather than the stability of particular solutions. Imposing structural stability restricts solution behaviour to only three possible types: stable (perturbations decay), unstable (perturbations grow), or saddle-like (mixed stability). That is, either there is nothing new, i.e., a perturbing solution rapidly dies out and the system returns to its unperturbed state, or diverges, ripping the system apart, or - as perhaps the worst alternative - a mixture of the two that results in an unstable saddle-like state. Consequently, such a system is in some sense always ‘trapped in itself’ being unable to pass to another, qualitatively different, state like for example one of those we observe around us (this will be made precise below). However, structural stability imposes very strong restrictions to the solutions describing different manifestations of the physical phenomena associated with these laws, for if a system is described by such equations then the set of its possible behaviours is restricted by the condition that any nearby system, i.e., any small perturbation of the system itself, behaves exactly as the original one [15]-[20].
1.4 Historical background
While this review builds on the author’s recent results [21][25], it is rooted in a longer tradition of research spanning four key areas of general relativity:
  • Linearization (in-)stability of the Einstein equations
  • Null hypersurfaces, caustics, wavefronts, and event horizons
  • Singularity theorems and the issue of the focusing state
  • Dynamical systems in cosmology and relativistic astrophysics
Below we briefly survey each area and highlight its connection to gravitational bifurcation theory. A common theme is the evolution problem of the Einstein equations and the physical plausibility of their solutions.
Early work on the linearization stability of the Einstein equations laid the foundation for understanding when small perturbations of a known solution again satisfy the full field equations [26]. Moncrief’s seminal analysis showed that, under suitable symmetry and boundary-condition assumptions, the space of nearby solutions forms a smooth manifold [27, 28]. Fischer and Marsden extended this by casting the Cauchy problem into a symmetric-hyperbolic framework and identifying constraint-equation obstructions that block linearized solutions from integrating to exact ones [29]-[31]. Since linearization stability of the Einstein equations is equivalent to that of their constraint equations [32], these works pinpointed when one can apply the implicit-function theorem to the constraint map [33].
The study of null hypersurfaces, caustics, and wavefronts dates back to Raychaudhuri’s and Sachs’s analysis of geodesic-congruence focusing [34]-[36], Penrose’s investigation of gravitational lensing and wavefront caustics [37], and Friedrich & Stewart’s Hamiltonian treatment of Legendrian and Lagrangian singularities in Minkowski space [38, 39]. These works revealed how singularities in the ‘optical’ equations signal breakdown of smooth null foliations, an early glimpse of dispersive behavior in the characteristic formulation.
The singularity theorems cemented the connection between causal focusing and inevitable spacetime singularities under very general energy and causality conditions [41]-[47]. Physically, once a congruence of null or timelike geodesics reaches a caustic, the classical evolution ‘breaks down’, just as a bifurcation in $\dot{x}=x^{2}$ signals a qualitative change in the solution set.
Finally, dynamical-systems methods in cosmology and relativistic astrophysics have roots in Misner’s mixmaster (Bianchi IX) model and the BKL oscillatory approach to singularities [48, 49], later elaborated by Barrow’s chaotic-cosmology work [50], and explored by Collins & Hawking [51, 52] and Bogoyavlenski & Novikov [53]. More recently, the monographs [3, 4, 40], together with Choptuik’s discovery of critical phenomena in gravitational collapse [54], have shown how ODE and PDE bifurcations underlie cosmic evolution and black-hole formation.
Together, these strands sketch general relativity as a theory rich in structural instabilities and pave the way for applying versal unfoldings and bifurcation theory techniques to classify every way in which symmetry-reduced gravitational systems can deform and bifurcate, the subject of Sections 2 and 3.
1.5 Bifurcation theory in gravitational context
Physically speaking, the opposite of structural stability can be said to lie with ‘bifurcating’ systems in which small changes can produce qualitatively new solutions. Bifurcation theory may be defined as the study of structurally unstable, also called ‘dispersive’, phenomena, and while its precise nature as a mathematical theory is probably unfamiliar to, or perhaps even might be considered as peripheral at best by, experts in the field of gravitation, the purpose of this paper is rather to suggest that there is a fundamental link between gravity and bifurcation theory and explore this relation using methods of the singularity theory of functions [^1] We shall use the words structural stability, hyperbolicity, and non-degeneracy as synonymous, and the word bifurcation as synonymous to structural instability, dispersive behaviour, non-hyperbolicity, and degeneracy. We shall also generally use ‘bifurcation theory’ as a generic phrase describing all three mathematical theories of ‘bifurcations’, ‘singularities’ (of functions), and ‘catastrophes’. For further discussion and proofs of the general bifurcation theory statements used below the reader is invited to consult Refs. [55]-[74], while for all statements related to gravitational bifurcations, we refer to [21]-[25]..
There is a great difference between structurally stable and dispersive systems of dynamical equations and for our present purposes a very characteristic manifestation of that difference is provided by the phenomenon of gravitational instability, as for example in the problem of cosmological structure formation where, while it usually treated in the linear approximation as in physical cosmology [75]-[85], the universe was very smooth but has become very inhomogeneous on small scales as cosmological observations clearly suggest. The universe ‘settles down’ to a novel state which albeit a changing one, it does not suddenly disappear but ‘persists’ for a time comparable to the observed lifetime of the system.
It is a standard geometric procedure in gravitational studies to symmetry-reduce the full Einstein field equations as applied to specialized situations describing phenomena of interest such as in cosmology, black holes, and other areas of gravity research. Most commonly, such a reduction is accompanied by the drastic transformation of the resulting differential systems from partial to ordinary differential equations, in other words reducing an infinite-dimensional degrees of freedom system to a finite-dimensional one, such as is the case for example in a Friedmann-Lemaítre context, or a Bianchi-type situation, or similar finitely-dimensional reductions in black hole spacetimes, etc. Such reductions often lead to interesting idealizations from a physical point of view and also to intriguing possibilities of adhering with observations. But while they lead to an obvious partial differential equation stability problem, that is whether the found symmetric ‘ordinary differential equation’ solution is stable or unstable in the full ‘unreduced’ context, they usually also reveal certain important features shared not only by the symmetrically reduced system but also by the original full (i.e., unsymmetric) system of partial differential equations: The symmetrically reduced systems obtained are typically structurally unstable. This is a common property not only met in the Einstein equations when applied to concrete symmetric equations, or analogous ones in modified gravity or string theories, but also in ‘optical’ type systems describing the evolution of congruences of causal curves or higher-dimensional gravitating objects in spacetime. The nature of the resulting structural instability of the symmetry-reduced laws can be succinctly summarized both physically and mathematically by the statement that:
@@ 1.5 Bifurcation theory in gravitational context
Furthermore, gravitational systems share two characteristic features which are in fact very akin to those of typical structurally unstable systems:
@@
    They contain parameters whose values are not really known exactly.
@@
    They are governed by differential equations which are highly sensitive to small modifications of their defining terms.
@@
The fluid parameter $\gamma$, or a cosmological constant are examples of the first property, while any ‘randomly chosen’ system of equations describing a gravity problem may be indicative of the second. Although a demand of structural stability is understandable as it implies that a property possessed by such a system is then physically very plausible, such a restriction imposed on an unstable system satisfying the two above conditions may lead to non-characteristic results attributed to the original equations, and leave out other behaviours which might lie closer to the nature of the system’s instabilities.
1.6 Key questions and scope of this review
@@
Another related issue is what does one really mean by a ‘stable solution’ in the context of a structurally unstable gravitational system? How may one classify the various unstable modes available in a given gravitational problem described and governed by a structurally unstable set of dynamical equations? Such issues point to a further consideration of related questions. For example, how can one perturb a structurally unstable gravitational system? What is the possible range of such perturbations, and how are the solutions of the original system being affected by the latter? What is the meaning of the phrase ‘small perturbation’ for a structurally unstable situation where there may be almost ignorable perturbations that have the ability to completely disrupt all orbits? Most importantly, how can one ‘lift’ all possible solution branches appearing at the level of the symmetry-reduced finitely-dimensional dynamical system to the full un-symmetric equation (say, the Einstein equations) and decide which one is finally picked?
The main message of this paper is that gravitation is a structurally unstable phenomenon and must be treated using an approach where this feature does not disappear altogether in the process but persists. In particular, the observed cosmic structures could have never been formed and persisted had gravity been structurally stable. How does then gravitational instability in its various forms manifest itself and related to the structural instability of the governing gravitational equations? How can one describe the latter in a mathematically precise yet physically consistent and plausible way? What are the main physical implications predicted by gravitational structural instability alone that could never have arisen had we treated the equations of a particular gravitational setup as being structurally stable?
@@
We shall return to some of these questions in the coming sections of this work, where we shall introduce the approach of bifurcation theory in the general framework of gravitational physics as a possible solution to this problem. Basically, very roughly speaking, this approach to the problem of general gravitational instabilities dictates that: one must first clearly recognise the kind or degree of degeneracy of the original system by bringing it to a ‘normal form’; secondly, embed the system of dynamical equations in a suitable larger parametric family customized to the given problem; and thirdly, prove that this family contains generically all possible stable perturbations of the original system. This is a long path, but eventually becomes very rewarding. One acquires control over all degeneracies of the original system, now in a structurally stable framework where each unstable mode persists, and in this way one is led to the complete spectrum of all possible behaviours associated with the original system.
1.7 Outline of the paper
@@
Section 2 gives a brief exposition of the main methods of bifurcation theory. In Section 2.1, we discuss the main differences between hyperbolic and nonhyperbolic systems, and in the remaining of Section 2 we detail on the main mathematical tools used in such an approach, with an eye to their subsequent application in a gravitational context (cf. Section 3). In Section 3, we study the picture which emerges from an application of these methods to some of the most basic and well-studied gravity problems in general relativity, cosmology, and black holes. In particular, in Section 3.1 we formulate the main problem addressed in this work and write down the central objectives in the search of gravitational bifurcations, while the remaining of this Section presents the main features of an application of bifurcation theory to certain important problems formulated after symmetry reduction of the Einstein equations. We also present the main results of such an analysis as applied to the Raychaudhuri equation and related systems, and Section 3.7 comprises a representative list of open problems associated with this approach. In the last Section, we discuss more general issues which emerge from gravitational bifurcations.
@@
While this work may be viewed as a review based on the results of Refs. [21]-[25], however, we present here a more qualitative way towards those results, with a view to the bigger picture which necessarily emerges and relates to a forming interface between general relativity and bifurcation theory. We consequently skip all mathematical details of the long arguments and calculations needed to prove many of the statements made in those references, but only referred to here in a ‘high-level’ manner. After reading the present paper, we hope that the reader will be sufficiently motivated to at least look at some of the proofs in those references.
2 Dynamics and bifurcations
In this Section, we describe very briefly some of the main methods in the vast territory of bifurcation theory in an effort to orient the reader in the maze of details and also to establish the notation we shall use subsequently in Section 3. We have developed the introductory material in the first few subsections, while later in this Section we shall discuss additional, somewhat more advanced, methodology topics.
Consider a set of dynamical equations that govern the evolution of a gravitating system, we shall call this set the original problem. For example, we imagine that this problem can be described mathematically by an autonomous dynamical system obtained from the Einstein equations after some symmetry idealization has been imposed, like with some of the problems analyzed in Section 3.
2.1 Hyperbolic vs. dispersive systems
2.1.1 General comments
@@
Besides the theory of small fluctuations mentioned in Section 1 in the context of physical cosmology, an essentially hyperbolic approach, i.e., assuming structural stability in the context of individual systems, is also used in other, somewhat more mathematical, treatments of gravity problems, cf. e.g., the pioneering works [2]-[6] and refs. therein, and has its roots in the standard theory of (hyperbolic) dynamical systems [15]-[20]. While this approach imposes two very important conditions on the gravity systems studied in this way, nevertheless it is the standard, well-known path that leads to systems with (usually complicated) hyperbolic equilibria: firstly, near such equilibria the equations become topologically equivalent to their linear parts, and secondly, ‘nearby’ nonlinear systems behave like their unperturbed counterparts.
In this work, we shall instead consider structurally unstable systems only, and follow their evolution as the parameters present in the problem vary. Our approach has a very high degree of applicability not only in cosmology but in other areas as well (for a description of possible cosmological areas of application of these ideas, cf. e.g., [7]), because - as we hope it will become clear in the following - the abundance of such systems in gravitation is high to such an extent as it may not be an exaggeration to say that a gravity problem obtained by a symmetry reduction of the Einstein (or ‘similar’) equations to a dynamical systems form is typically structurally unstable.
@@
To motivate gravitational bifurcation theory employed in this work, let us recall the ways in which difference between a hyperbolic and a non-hyperbolic system arises (a hyperbolic system is one whose all equilibria, i.e., steady state solutions, are hyperbolic). Equilibria of a hyperbolic system are always isolated and the behaviour of solutions near them is qualitatively the same as that of their linear parts. However, dispersive systems may exhibit bifurcating behaviour and as a rule comprise multiple solutions intertwined with dramatic changes in the behaviour of the orbits. In addition, their global dynamics may be fully described by ‘restricting’ to only those dimensions which correspond to the non-hyperbolic eigenvalues (i.e., tangent to the null eigenspaces).
2.1.2 Hyperbolic dynamics